A student buys a bag of jelly beans at the store. She eats most of them, but 20 remain at the end of the day. Exactly 13 of these jelly beans are fruity, and the other 7 are root beer. Her boyfriend grabs 5 of the 20 jelly beans.

a.) what is the probability that exactly 3 out of the 5 jelly beans that he grabs are fruity?

b.) what is the probability that the boyfriend grabs at least 1 root beer jelly bean?
Answer #1

a)

probability that exactly 3 out of the 5 jelly beans that he grabs are fruity =P(select 3 of 13 fruity and 2 from 7 root beer)= =286*21/15504 =0.387384

b)

probability that the boyfriend grabs at least 1 root beer jelly bean =1-P(no root beer jelly bean )

=1- =1-1287*1/15504=1-0.083011=0.916989